# Analog Quantum Simulation of Coupled Electron-Nuclear Dynamics in Molecules

### v1

@@ II.1 Analog quantum simulation
We now introduce our approach to map the Hamiltonian $\hat{H}_{\mathrm{mol}}$ onto an analog quantum device to perform a pre-BO simulation. Such a mapping must satisfy the symmetry requirements of fermions (electrons) and bosons (vibrations), given by their (anti-)commutation relations. Ideally, we would like to map both types of degrees of freedom onto a device with native fermionic and bosonic degrees of freedom. There are quantum architectures that contain bosonic modes such as ion traps and circuit quantum electrodynmics (cQED) devices, and some basic architectures with controllable fermionic states [47, 48] have been proposed. However, there is currently no practical quantum architecture that contains both degrees of freedom with individual control. Instead, we will use a fermion-qubit mapping to represent spin orbital ladder operators with non-local qubit operators [49, 50]. This allows us to use ion traps or cQED, since both have qubit levels in addition to their bosonic modes. We will refer to architectures that contain multiple qubits and bosonic modes with controllable coupling between them as “coupled multi-qubit-boson” (cMQB) devices, which are a subset of mixed qudit-boson (MQB) devices [39].
@@
An illustration of our pre-BO analog simulation method in comparison with the BO framework is summarized in Fig. 1. After finding a nuclear position dependent spin orbital basis (Fig. 1a) and its expansion coefficients in Eq. 3, the pre-BO wavefunction (Fig. 1b) is encoded on the cMQB simulator, e.g. on a trapped-ion device with ion electronic states representing electronic ONVs, and ion motional modes representing the nuclear component of the wavefunction (Fig. 1c). In contrast, existing classical/quantum algorithms in the BO framework require the pre-calculation of BO electronic states (Fig. 1d) with the spin orbital basis. The vibronic wavefunction is then propagated on a truncated BO basis (Fig. 1e), which can result in inaccurate dynamics simulations.
II.1.1 Fermion-qubit mapping
@@
where $N_{\mathrm{q}}$ is the number of qubits, $\hat{P}^{I}_{k}$ is one of the identity or Pauli operators for the $k$-th qubit ($\hat{P}_{k}\in\{\hat{I}_{k},\hat{X}_{k},\hat{Y}_{k},\hat{Z}_{k}\}$) in the $I$-th Pauli string, and $f_{I}$ is a function of bosonic ladder operators $\hat{b}_{\nu}$ and $\hat{b}_{\nu}^{\dagger}$ coupled to the $I$-th Pauli string. Although the fermion-qubit mapping and the corresponding electronic basis is not spin-adapted, the Hamiltonian commutes with total electronic spin $z$-projection and square ($\hat{S}_{z}$ and $\hat{S}^{2}$) operators. Correspondingly, the electronic part of the molecular vibronic state $|\mathbf{n}\rangle_{\mathrm{e}}$ in our ansatz (Eq. 2) is mapped to the multi-qubit state $|\mathbf{q}\rangle_{\mathrm{q}}=|q_{1}\cdots q_{N_{\mathrm{q}}}\rangle_{\mathrm{q}}$ of the cMQB device, while the multi-mode vibrational wavefunction is directly mapped to the bosonic degrees of freedom (Fig. 1c). A schematic illustration of a circuit for pre-BO dynamics simulation with the cMQB mapping consists of initialization, time-evolution, and measurement is shown in Fig. 2.
II.1.2 Initialization
@@
As with any quantum dynamics simulation, our approach requires a robust preparation of the initial state which can be generated by a unitary operator $\hat{F}$ (Fig. 2a). For most applications in photochemical dynamics we can assume that the initial state is well represented as a product state of electronic and nuclear components of the wavefunction which can be prepared by $\hat{F}=\hat{F}_{\mathrm{e}}\otimes\hat{F}_{\mathrm{n}}$, with separate operations on qubits with $\hat{F}_{\mathrm{e}}$ and motional states with $\hat{F}_{\mathrm{n}}$. According to the Franck-Condon approximation, the equilibrium nuclear state is unperturbed by photoexcitation and the excitation is instantaneous. This means the initial state can be prepared with a ground (or coherent) nuclear state, and with qubit states corresponding to the photoexcited spin orbital occupation.
@@
Future devices with longer coherence times and more sophisticated quantum control will allow for increasingly accurate initial state preparation. For example, controlled dissipation of the simulator could be used to prepare the vibronic ground state wavefunction, with the fidelity approaching 1 for longer experimental times. The dynamics could then be initialized by a purely electronic (Franck-Condon) or vibronic transition dipole operator. By implementing a time-dependent Hamiltonian with the transition dipole operator, our approach can also serve as a platform for the explicit simulation of vibronic effects in light-matter interactions.
II.1.3 Coupled multi-qubit-boson time-evolution operator
@@
The direct implementation of the approach outlined above requires the simultaneous control of all degrees of freedom in the simulation. This becomes impractical for large molecules with many vibrational modes. Therefore, instead of simultaneous evolution, we can use Trotterization [51] of the time-evolution operator. Trotterization approximates the time-evolution operator as a product of operators, each of which corresponds to a term $\hat{H}_{I}$ in Eq. 4. The Trotter time-evolution operator is given by $\hat{U}_{\mathrm{cMQB}}(t)\approx\left(\prod^{N_{\mathrm{op}}}_{I}\hat{U}_{I}(t/N_{\mathrm{t}})\right)^{N_{\mathrm{t}}}\equiv\hat{U}(\Delta t)^{N_{\mathrm{t}}}$, where $N_{\mathrm{op}}$ is the number of divided operators and $N_{t}$ is the number of the Trotter steps (Fig. 2b).
@@
The time-evolution operator for each term in our Hamiltonian (Eq. 4) is given by
@@
where $\hat{\Lambda}^{I}$ is a product of tensor products of Clifford gates (Hadamard, phase, and CNOT gates) to propagate the coupling between bosonic mode(s) and a single qubit $q_{0}^{I}$ generated by a laser-ion interaction to multiple qubits for the desired Pauli strings for $\hat{H}_{I}$. In Fig. 2b, we show an example cMQB time-evolution operator $\hat{U}_{I}(\Delta t)=\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{Y}_{2}\hat{Z}_{3}\hat{Y}_{4})$, which can be achieved by a sequence of digital and analog quantum operations: $\hat{\mathrm{S}}_{4}\hat{\mathrm{S}}_{2}\hat{\mathrm{H}}_{3}\hat{\mathrm{C}}_{34}\hat{\mathrm{C}}_{32}\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{X}_{3})\hat{\mathrm{C}}_{32}\hat{\mathrm{C}}_{34}\hat{\mathrm{H}}_{3}\hat{\mathrm{S}}_{2}^{\dagger}\hat{\mathrm{S}}_{4}^{\dagger}$ where $\hat{\mathrm{C}}_{ij}$ is the CNOT gate for the control qubit $i$ and the target qubit $j$, and $\hat{\mathrm{H}}_{i}$ and $\hat{\mathrm{S}}_{i}$ are the Hadamard and phase gates for qubit $i$, respectively. Therefore, this approach can be considered a digital-analog quantum simulation [52] because the gates involved in generating qubit entanglement take a digital form. A fully analog form of the time-evolution could be achieved using time- and spin-dependent squeezing/displacement operators on the collective ionic motions [53]; however, for the remainder of this paper we consider the digital-analog approach.
@@
Figure 2: A schematic quantum circuit diagram for pre-BO vibronic dynamics simulation with the cMQB approach. Straight lines represent qubits and wavy lines represent the motional modes. a, A simulation consists of initialization, time-evolution, and measurement stages. An initial vibronic state can be prepared by a unitary operator $\hat{F}$. The initial state then propagated by the time-evolution operator $\hat{U}(t)$. Finally, an expectation value can be measured with Hadamard tests with the corresponding operator $\hat{O}$ controlled by an ancilla qubit state (represented by a black circle) using Hadamard ($\hat{\mathrm{H}}$) and phase gates ($\hat{\mathrm{S}}$). Real and imaginary parts of the expectation value are obtained with and without the phase gate, respectively. b, Trotterization of the time-evolution operator $\hat{U}(t)$ and the digital-analog circuit for an example Trotterized cMQB operator $\hat{U}(\Delta t)=\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{Y}_{2}\hat{Z}_{3}\hat{Y}_{4})$. CNOT gates are represented by symbols with a black circle for the control qubit, connected to an open circle for the target qubit.
@@
cMQB devices such as trapped ions and cQED typically use bosonic modes to achieve entangling gates in digital quantum algorithms. The quantum gates necessary to implement $\hat{\Lambda}^{I}$ are thus readily achieved with these architectures. However, because the entangling gates require a bosonic mode, certain bosonic modes must be reserved from those used to simulate nuclear vibrational degrees of freedom. The total number of bosonic modes in the simulator must therefore be at least one greater than the total number of molecular vibrational modes.
II.1.4 Measurement
@@
Our approach encodes the electronic degrees of freedom of the vibronic wavefunction as an ONV, and nuclear degrees of freedom are mapped to the bosonic modes. This difference in encoding means that all nuclear observables can be measured directly from the simulator, whereas some electronic and vibronic observables will depend on the electronic orbital basis functions. In general, the expectation values of observables can be obtained with Hadamard tests [54] for the corresponding unitary operator $\hat{O}$, in which an ancilla qubit in a superposition of $|0\rangle$ and $|1\rangle$ states is used to control the operator on the wavefunction (Fig. 2a). Measurement of the ancilla in the $\hat{X}$ basis yields the real component of the expectation value of the operator, and the $\hat{Y}$ basis measurement gives the imaginary component.
@@
$\rho(\mathbf{r},\mathbf{Q},t)$ $=\sum_{p,q}\Bigg{[}\varphi_{p}^{*}(\mathbf{r};\mathbf{Q})\varphi_{q}(\mathbf{r};\mathbf{Q})\delta_{\sigma_{p}\sigma_{q}}$
@@
II.2 Connection to the Born-Oppenheimer framework
@@
If we define the electronic Hamiltonian $\hat{H}^{\mathrm{e}}(\mathbf{Q})$ as the first two terms of Eq. 3 at a fixed nuclear position $\mathbf{Q}$, then we can find an electronic Hamiltonian matrix $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ with elements $H_{\mathbf{mn}}^{\mathrm{e}}(\mathbf{Q})=\langle\mathbf{m}|\hat{H}^{\mathrm{e}}(\mathbf{Q})|\mathbf{n}\rangle_{\mathrm{e}}$, where only electronic degrees of freedom are integrated. The Hamiltonian matrix can be transformed by a unitary matrix $\mathbf{W}(\mathbf{Q})$ to give $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})=\mathbf{W}(\mathbf{Q})\mathbf{H}^{\mathrm{e}}(\mathbf{Q})\mathbf{W}^{\dagger}(\mathbf{Q})$. When $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})$ is a fully diagonal matrix, we arrive at the exact (full configuration interaction, FCI) solution to the electronic structure problem. In practice, full diagonalization is intractable on classical computers, so the eigenvalues of a submatrix of $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ may be found instead. Iterative diagonalization is typically used to further reduce the cost, and the number of eigenvalues found, $N_{\mathrm{BO}}$, can be less than the rank of the (sub)matrix. The BO electronic states are the corresponding eigenfunctions (Fig. 1d), and molecular dynamics can be expressed by the Born-Huang expansion as nuclear dynamics on multiple BO states:
@@
where $\mathbf{n}({\underline{\mathbf{x}}};\mathbf{Q})=\langle{\underline{\mathbf{x}}},\mathbf{Q}|\mathbf{n}\rangle$ and $\mathbf{v}(\mathbf{Q})=\langle\mathbf{Q}|\mathbf{v}\rangle$ are the real space wavefunctions of the electronic and nuclear ONVs, i.e. the Slater determinant and Hartree product of single particle basis functions, respectively. $W_{j\mathbf{n}}$ is an element of $\mathbf{W}$, $\Phi_{j}^{\mathrm{BO}}({\underline{\mathbf{x}}};\mathbf{Q})=\sum_{\mathbf{n}}W_{j\mathbf{n}}\mathbf{n}({\underline{\mathbf{x}}};\mathbf{Q})$ is the $j$-th BO state, and $\chi^{\mathrm{BO}}_{j}(\mathbf{Q},t)=\sum_{\mathbf{v}}T_{j\mathbf{v}}(t)\mathbf{v}(\mathbf{Q})$ is the corresponding time-dependent BO-projected nuclear wavefunction with an expansion coefficient $T_{j\mathbf{v}}(t)$ (Fig. 1e).
@@
Without spin-orbit coupling, the electronic Hamiltonian commutes with the electron number operator, and the $z$-component and square of the total electron spin angular momentum operator. The maximum number of BO electronic states is thus equal to the number of configuration state functions (CSFs, i.e. spin-adapted linear combinations of Slater determinants) with a fixed electron count and spin, $N_{\mathrm{CSF}}$, which we will discuss in more detail in Sec IV. The summation over the BO states in Eq. 8 is exact within the given orbital basis set only when $N_{\mathrm{BO}}=N_{\mathrm{CSF}}$. The relation between the Born-Huang expansion and our ansatz (Eq. 2) becomes clear via the relation $\sum_{j}^{N_{\mathrm{CSF}}}\sum_{\mathbf{v}^{\prime}}T_{j\mathbf{v}^{\prime}}(t)(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=C_{\mathbf{v}\mathbf{n}}(t)$, where $(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=\int\mathrm{d}\mathbf{Q}\,\mathbf{v}^{*}(\mathbf{Q})W_{j\mathbf{n}}(\mathbf{Q})\mathbf{v}^{\prime}(\mathbf{Q})$. However, $N_{\mathrm{CSF}}$ scales rapidly with the orbital basis set size whereas $N_{\mathrm{BO}}$ must be kept small for practical simulations. Therefore, the GBOA is employed ($N_{\mathrm{BO}}<N_{\mathrm{CSF}}$) to reduce the computational cost of the BO state calculations and time-propagation of the molecular vibronic state (Fig. 1e). In contrast, our ansatz (Eq. 2) is exact within the given vibronic basis by definition (Fig. 1c).
III Numerical test

### v2

@@ II.2 Analog quantum simulation
We now introduce our approach to map the Hamiltonian $\hat{H}_{\mathrm{mol}}$ onto an analog quantum device to perform a pre-BO simulation. Such a mapping must satisfy the symmetry requirements of fermions (electrons) and bosons (vibrations), given by their (anti-)commutation relations. Ideally, we would like to map both types of degrees of freedom onto a device with native fermionic [50, 51] and bosonic degrees of freedom. However, there is currently no practical quantum architecture that contains both degrees of freedom with individual control. Therefore, we will target quantum architectures that have qubit levels and bosonic modes with controllable coupling between them, such as ion traps and circuit quantum electrodynamics (cQED), using a fermion-qubit mapping [52, 53]. We will refer to such architectures as “coupled multi-qubit-boson” (cMQB) devices.
@@
An illustration of our pre-BO analog simulation method in comparison with the BO framework is summarized in Fig. 1. We start by finding a nuclear position dependent spin orbital basis [Fig. 1(a)] and deriving the expansion coefficients in Eq. (3). We then encode the pre-BO wavefunction [Fig. 1(b)] on a cMQB device, e.g. a trapped-ion device with ion electronic states representing electronic ONVs, and ion motional modes representing the nuclear component of the wavefunction [Fig. 1(c)]. In contrast, existing classical/quantum algorithms in the BO framework require the pre-calculation of BO electronic states [Fig. 1(d)] with the spin orbital basis. The vibronic wavefunction is then propagated on a truncated BO basis [Fig. 1(e)], which can result in inaccurate dynamics simulations.
@@
where $N_{\mathrm{q}}$ is the number of qubits, $\hat{P}^{I}_{k}$ is one of the identity or Pauli operators for the $k$-th qubit ($\hat{P}_{k}\in\{\hat{I}_{k},\hat{X}_{k},\hat{Y}_{k},\hat{Z}_{k}\}$) in the $I$-th Pauli string, and $f_{I}$ is a function of bosonic ladder operators $\hat{b}_{\nu}$ and $\hat{b}_{\nu}^{\dagger}$ coupled to the $I$-th Pauli string. Correspondingly, the electronic part of the molecular vibronic state $|\mathbf{n}\rangle_{\mathrm{e}}$ in our ansatz [Eq. (2)] is mapped to the multi-qubit state $|\mathbf{q}\rangle_{\mathrm{q}}=|q_{1}\cdots q_{N_{\mathrm{q}}}\rangle_{\mathrm{q}}$ of the device, while the multi-mode vibrational wavefunction is directly mapped to the bosonic degrees of freedom [Fig. 1(c)]. The molecular cMQB Hamiltonian may be scaled with a factor $F$ to a simulation scale ($\hat{H}_{\mathrm{sim}}=F\hat{H}_{\mathrm{mol}}$) which depends on the experimental parameters of the simulator.
Our electronic basis spans the entire electronic Fock space, including electron configurations with different number of electrons and spins. This implies that our approach can describe intersystem crossing if the corresponding coefficients derived from integrals of the spatial orbitals and electron angular momentum are non-zero. Our approach is also capable of describing ionization processes by adding interactions with terms that remove electrons, which has potential uses for predicting photoelectron spectra. In both cases, the number of quantum resources is unchanged for the simulation.
A schematic illustration of a circuit for pre-BO dynamics simulation with the cMQB mapping consists of initialization, time-evolution, and measurement, as shown in Fig. 2. We provide details for each step in what follows.
@@
As with any quantum dynamics simulation, our approach requires a robust preparation of the initial state which can be generated by a unitary operator $\hat{U}_{\mathrm{init}}$ [Fig. 2(a)]. For most applications in photochemical dynamics we can assume that the initial state is well represented as a product state of electronic and nuclear components of the wavefunction which can be prepared by $\hat{U}_{\mathrm{init}}=\hat{U}_{\mathrm{e}}\otimes\hat{U}_{\mathrm{n}}$, with separate operations on qubits with $\hat{U}_{\mathrm{e}}$ and motional states with $\hat{U}_{\mathrm{n}}$. According to the Franck-Condon approximation, the equilibrium nuclear state is unperturbed by photoexcitation and the excitation is instantaneous. This means the initial state can be prepared with a ground (or coherent) nuclear state, and with qubit states corresponding to the photoexcited spin orbital occupation.
@@
Future devices with longer coherence times and more sophisticated quantum control will allow for increasingly accurate initial state preparation. For example, controlled dissipation of the simulator could be used to prepare the vibronic ground state wavefunction, with the fidelity approaching 1 for longer experimental times. The dynamics could then be initialized by a purely electronic (Franck-Condon) or vibronic transition dipole operator.
@@
The direct implementation of our approach outlined above requires the simultaneous control of all degrees of freedom in the simulation. However, our approach involves a fermion-qubit mapping on the electronic degrees of freedom, which involves a change of basis between non-commuting electronic terms. Therefore, we must use Trotterization [54] of the time-evolution operator. Trotterization approximates the time-evolution operator as a product of operators, each of which corresponds to a term $\hat{H}_{I}$ in Eq. (4). The Trotterized cMQB time-evolution operator is given by $\hat{U}(t)\approx\left(\prod^{N_{\mathrm{op}}}_{I}\hat{U}_{I}(t/N_{\mathrm{t}})\right)^{N_{\mathrm{t}}}\equiv\hat{U}_{\mathrm{cMQB}}(\Delta t)^{N_{\mathrm{t}}}$, where $N_{\mathrm{op}}$ is the number of divided operators and $N_{\mathrm{t}}$ is the number of the Trotter steps [Fig. 2(b)].
@@
The Trotterized cMQB time-evolution operator for each term with an electronic component in our Hamiltonian [Eq. (4)] is given by
@@
where $\hat{\Lambda}^{I}$ is an $N$-qubit Clifford operation which can be composed out of Hadamard, phase, and CNOT gates, to propagate the coupling between bosonic mode(s) and a single qubit $q_{0}^{I}$ generated by a laser-ion interaction to multiple qubits for the desired Pauli strings for $\hat{H}_{I}$. In Fig. 2(b) we show an example time-evolution operator $\hat{U}_{I}(\Delta t)=\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{Y}_{2}\hat{Z}_{3}\hat{Y}_{4})$, which can be achieved by a sequence of digital and analog quantum operations. The nuclear-only terms of the Hamiltonian (the final two terms of Eq. (3) are given up to second order by a constant detuning from the bosonic mode frequency. Higher-order nuclear-only terms can be achieved with operations on only the bosonic modes, or qubit-boson operations with an ancilla qubit. This approach can be considered a digital-analog quantum simulation [55, 56] because the gates involved in generating qubit entanglement take a digital form. A fully analog form of the time-evolution could be achieved using time- and spin-dependent squeezing/displacement operators on the collective ionic motions [57]; however, for the remainder of this paper we consider the digital-analog approach.
@@
Figure 2: A schematic quantum circuit diagram for pre-BO vibronic dynamics simulation with our cMQB approach. Straight lines represent qubits and wavy lines represent the motional modes. (a) A simulation consists of initialization, time-evolution, and measurement stages. An initial vibronic state can be prepared by a unitary operator $\hat{U}_{\mathrm{init}}$. The initial state is then propagated by the time-evolution operator $\hat{U}(t)$. Finally, an expectation value can be measured with Hadamard tests with the corresponding operator $\hat{O}$ controlled by an ancilla qubit state (represented by a black circle) using Hadamard ($\hat{\mathrm{H}}$) and phase gates ($\hat{\mathrm{S}}$). Real and imaginary parts of the expectation value are obtained with and without the phase gate, respectively. (b) Trotterization of the time-evolution operator $\hat{U}(t)$ and the digital-analog circuit for an example Trotterized operator $\hat{U}_{I}(\Delta t)=\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{Y}_{2}\hat{Z}_{3}\hat{Y}_{4})$. CNOT gates are represented by symbols with a black circle for the control qubit, connected to an open circle for the target qubit.
@@
Trapped ions and cQED typically use bosonic modes to achieve entangling gates in digital quantum algorithms. The quantum gates necessary to implement $\hat{\Lambda}^{I}$ are thus readily achieved with these architectures. However, since the entangling gates require a bosonic mode, some bosonic modes must be reserved from those used for simulating nuclear vibrational degrees of freedom. We refer to these as the bus modes. The total number of bosonic modes in the simulator must therefore be at least one greater than the total number of molecular vibrational modes.
@@
Our approach encodes the electronic degrees of freedom of the vibronic wavefunction as an ONV, and nuclear degrees of freedom are mapped to the bosonic modes. This difference in encoding means that all nuclear observables can be measured directly from the simulator, whereas some electronic and vibronic observables will depend on the electronic orbital basis functions. In general, the expectation values of observables can be obtained with Hadamard tests [58] for the corresponding unitary operator $\hat{O}$ as shown in Fig. 2(a).
@@
As an analog simulation approach, our approach is subject to environmental noise over the course of the simulation. The noise effect on an open quantum system can be described by the Lindblad master equation [60]:
$\dfrac{d\hat{\rho}}{dt}=-i[\hat{H},\hat{\rho}]+\sum_{i}\gamma_{i}D[\hat{L}_{i}]\hat{\rho},$ (8)
where $\hat{\rho}=|\Psi\rangle\langle\Psi|$ represents the density operator of the system, $\hat{H}$ is the Hamiltonian of the system, $D[\hat{L}]\hat{\rho}=\hat{L}\hat{\rho}\hat{L}^{\dagger}-\frac{1}{2}\{\hat{L}^{\dagger}\hat{L},\hat{\rho}\}$ is the Lindblad superoperator, $\hat{L}_{i}$ is the jump operator for the noise source $i$, and $\gamma_{i}$ is the corresponding dissipation rate. The noise effect on a cMQB device translates into the molecular state during a cMQB simulation, where the native rates of the cMQB simulator are scaled to the molecular system depending on the encoding, device parameters, and simulation setup.
The noise does not equally affect electronic and vibrational degrees of freedom in cMQB simulations, since they are encoded in different resources of the simulator. In addition, because cMQB simulators use motional modes to achieve entangling gates, the electronic degrees of freedom are subject to noise on both qubit states and motional modes.
In ion-trap cMQB devices, the dominant source of noise is the decoherence of the motional modes, with a rate orders of magnitude greater than the other sources of noise [61]. Therefore, in what follows, we only consider the noise effect from motional decoherence ($\hat{L}_{\nu}=\hat{b}_{\nu}^{\dagger}\hat{b}_{\nu}$) in an ion-trap cMQB device. While the motional modes representing the nuclear degrees of freedom are subject to the decoherence for the entire experiment time, each qubit is subject to the motional noise only when an entanglement operator is applied on the qubit. Thus, the native rates for the motional decoherence and the indirect qubit noise scale differently for vibrational and electronic degrees of freedom of the molecular state. The native decoherence rate $\gamma^{\mathrm{d}}_{\mathrm{mot}}$ of motional modes for a cMQB device is scaled to the molecular vibrational decoherence rate $\gamma^{\mathrm{vib}}_{\mathrm{mol}}$ as
$\gamma^{\mathrm{vib}}_{\mathrm{mol}}=\left(N_{\mathrm{CNOT}}\dfrac{t_{\mathrm{CNOT}}}{\Delta t_{\mathrm{mol}}}+\dfrac{N_{\mathrm{op}}}{F}\right)\gamma^{\mathrm{d}}_{\mathrm{mot}},$ (9)
while the average indirect spin noise ($\hat{L}_{q}\approx\hat{S}_{q}$) for each qubit in the molecular scale $\gamma_{\mathrm{mol}}^{\mathrm{q}}$ can be connected to the native motional decoherence rate of the bus mode:
$\gamma^{\mathrm{q}}_{\mathrm{mol}}=\dfrac{N_{\mathrm{CNOT}}}{4N_{\mathrm{o}}}\dfrac{t_{\mathrm{CNOT}}}{\Delta t_{\mathrm{mol}}}\gamma^{\mathrm{d}}_{\mathrm{mot}},$ (10)
where $\Delta t_{\mathrm{mol}}$ is the Trotter step size in the molecular scale, $t_{\mathrm{CNOT}}$ is the experimental time for a two-qubit entangling gate, and $N_{\mathrm{CNOT}}$ is the number of two-qubit entanglement gates per single Trotter step (See Supplemental Material [46]). Although a small Trotter step $\Delta t_{\mathrm{mol}}$ reduces the theoretical error of the cMQB time-evolution operator, the above equations imply that trade-off exists due to the number of digital qubit entanglement operations.
II.3 Connection to the Born-Oppenheimer framework
@@
If we define the electronic Hamiltonian $\hat{H}^{\mathrm{e}}(\mathbf{Q})$ as the first two terms of Eq. (3) at a fixed nuclear position $\mathbf{Q}$, then we can find an electronic Hamiltonian matrix $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ with elements $H_{\mathbf{mn}}^{\mathrm{e}}(\mathbf{Q})=\langle\mathbf{m}|\hat{H}^{\mathrm{e}}(\mathbf{Q})|\mathbf{n}\rangle_{\mathrm{e}}$, where only electronic degrees of freedom are integrated. The Hamiltonian matrix can be transformed by a unitary matrix $\mathbf{W}(\mathbf{Q})$ to give $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})=\mathbf{W}(\mathbf{Q})\mathbf{H}^{\mathrm{e}}(\mathbf{Q})\mathbf{W}^{\dagger}(\mathbf{Q})$. When $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})$ is a fully diagonal matrix, we arrive at the exact (full configuration interaction, FCI) solution to the electronic structure problem. In practice, full diagonalization is intractable on classical computers, so the eigenvalues of a submatrix of $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ may be found instead. Iterative diagonalization is typically used to further reduce the cost, and the number of eigenvalues found, $N_{\mathrm{BO}}$, can be less than the rank of the (sub)matrix. The BO electronic states are the corresponding eigenfunctions [Fig. 1(d)], and molecular dynamics can be expressed by the Born-Huang expansion as nuclear dynamics on multiple BO states:
@@
where $\mathbf{n}({\underline{\mathbf{x}}};\mathbf{Q})=\langle{\underline{\mathbf{x}}},\mathbf{Q}|\mathbf{n}\rangle$ and $\mathbf{v}(\mathbf{Q})=\langle\mathbf{Q}|\mathbf{v}\rangle$ are the real space wavefunctions of the electronic and nuclear ONVs, i.e. the Slater determinant and Hartree product of single particle basis functions, respectively. $W_{j\mathbf{n}}$ is an element of $\mathbf{W}$, $\Phi_{j}^{\mathrm{BO}}({\underline{\mathbf{x}}};\mathbf{Q})=\sum_{\mathbf{n}}W_{j\mathbf{n}}\mathbf{n}({\underline{\mathbf{x}}};\mathbf{Q})$ is the $j$-th BO state, and $\chi^{\mathrm{BO}}_{j}(\mathbf{Q},t)=\sum_{\mathbf{v}}T_{j\mathbf{v}}(t)\mathbf{v}(\mathbf{Q})$ is the corresponding time-dependent BO-projected nuclear wavefunction with an expansion coefficient $T_{j\mathbf{v}}(t)$ [Fig. 1(e)].
@@
Without spin-orbit coupling, the maximum number of BO electronic states is equal to the number of configuration state functions (CSFs, i.e. spin-adapted linear combinations of Slater determinants) with a fixed electron count and spin, $N_{\mathrm{CSF}}$, which we will discuss in more detail in Sec III. The summation over the BO states in Eq. (11) is exact within the given orbital basis set only when $N_{\mathrm{BO}}=N_{\mathrm{CSF}}$. The relation between the Born-Huang expansion and our ansatz [Eq. (2)] becomes clear via the relation $\sum_{j}^{N_{\mathrm{CSF}}}\sum_{\mathbf{v}^{\prime}}T_{j\mathbf{v}^{\prime}}(t)(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=C_{\mathbf{v}\mathbf{n}}(t)$, where $(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=\int\mathrm{d}\mathbf{Q}\,\mathbf{v}^{*}(\mathbf{Q})W_{j\mathbf{n}}(\mathbf{Q})\mathbf{v}^{\prime}(\mathbf{Q})$. However, $N_{\mathrm{CSF}}$ scales rapidly with the orbital basis set size whereas $N_{\mathrm{BO}}$ must be kept small for practical simulations. Therefore, the GBOA is employed ($N_{\mathrm{BO}}<N_{\mathrm{CSF}}$) to reduce the computational cost of the BO state calculations and time-propagation of the molecular vibronic state [Fig. 1(e)]. In contrast, our ansatz [Eq. (2)] is exact within the given vibronic basis by definition.
II.4 Numerical test